OpenStudy (anonymous):

sqrt 7x (sqrt x -7 sqrt 7) ?

OpenStudy (anonymous):

\[\sqrt{7x}(\sqrt{x}-7\sqrt{7})\]

OpenStudy (anonymous):

@$w3G_Godd

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

\[\sqrt{a} * \sqrt{b} = \sqrt{a*b}\]

OpenStudy (anonymous):

So first use distributive property that is if we have x * (a - b) = x * a - x * b Then use above rule ,,

OpenStudy (anonymous):

So using distributive property we get \[\sqrt{7x}*(\sqrt{x } - 7\sqrt{7}) = \sqrt{7x}*\sqrt{x} - \sqrt{7x}*7\sqrt{7} \]

OpenStudy (anonymous):

Now use √a∗√b=√(a*b) So we get \[\sqrt{7x^2} - 7*\sqrt{x*7^2} = x \sqrt{7} - 7*7\sqrt{x}= x \sqrt{7} - 49\sqrt{x}\]

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

You are welcome :) Sam :)

OpenStudy (anonymous):

sorry i just saw this :(